Prove that the product of any three consecutive integers is a multiple of 3.ย 


1
Expert's answer
2021-11-30T18:10:17-0500

Let the integers be "n, n + 1" and "n+2, n\in \Z."

If "n" is divisible by "3," then the product "n(n + 1)(n + 2)" is divisible by "3."

If "n" is not divisible by "3," then either "n = 3l + 1, l\in \Z" or "n = 3m + 2,"

"m\in \Z."

By the division algorithm

In the first case, "n + 2 = 3l + 3 = 3(l + 1), l\in \Z" is divisible by 3 and therefore the product


"n(n + 1)(n + 2) = 3(3l+1)(3l + 2)(l+1)"

is divisible by 3.

In the second case, "n + 1 = 3m +3 = 3(m + 1), m\in \Z" is divisible by 3 and therefore the product


"n(n + 1)(n + 2) = 3(3m+2)(m+1)(3m+4)"

is divisible by 3.

This shows that the product of any three consecutive integers is a multiple of "3."